# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v2

@@ Abstract
Conventional survival metrics, such as Harrell’s concordance index and the Brier Score, rely on the independent censoring assumption for valid inference in the presence of right-censored data. However, when instances are censored for reasons related to the event of interest, this assumption no longer holds, as this kind of dependent censoring biases the marginal survival estimates of popular nonparametric estimators. In this paper, we propose three copula-based metrics to evaluate survival models in the presence of dependent censoring, and design a framework to create realistic, semi-synthetic datasets with dependent censoring to facilitate the evaluation of the metrics. Our empirical analyses in synthetic and semi-synthetic datasets show that our metrics can give error estimates that are closer to the true error, mainly in terms of prediction accuracy.
@@ 1 Introduction
Learning survival prediction models resembles learning regression models, as both are essentially trying to learn, from a labeled dataset, a model that maps a description of an instance to a real value. However, survival prediction differs as some of the training instances are “censored” – that is, we only know a lower bound of the true time. This includes, for example, individuals who are still alive at the end of the study, as well as others who left the study and were “lost to follow-up”. The independent censoring assumption states that, within any subgroup of interest, we assume that the patients (i.e., instances) who are censored at time $t$ are representative of all the patients in that subgroup who remained at risk at time $t$ with respect to their survival experience [Emura and Chen, 2018, Ch. 1]. That is, censoring is random within the subgroups. Let $E$ be the event distribution and $C$ be the censoring distribution, then we have $E\perp C$i.e.$E$ provides no information about the distribution of ${C}$, and vice versa [Kleinbaum and Klein, 2012, Ch. 1]. If there are features (covariates) $\bm{X}$, this assumption becomes the conditional independent censoring assumption, i.e. $E\perp C\mid\bm{X}$ [Kleinbaum and Klein, 2012, Ch. 1]. (See Appendix C for details.) Note that many standard survival tools – e.g., the Kaplan-Meier (KM) [Kaplan and Meier, 1958] and Nelson-Aalen [Nelson, 1969, Aalen, 1978] estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [Cox, 1972], rely on an assumption of independent censoring for valid inference in the presence of right-censored data [Kleinbaum and Klein, 2012, Ch. 1].
@@
Figure 1: (a) Residual dependence between event time $E$ (cancer relapse) and censoring time $C$ (cancer death) denoted by the diamond ‘Dep’, due to an unobserved confounding covariate – here ‘Tumor grade’. (b) Evaluation under dependent censoring, where Patient A is censored immediately before the event for reasons related to the event. This leads to a bias estimation on the marginal survival distribution using the Kaplan-Meier (KM) [Kaplan and Meier, 1958] estimator. Note that the two curves are coincide between range $[0,c_{i}]$. (c) Scatter plot generated from a standard exponential distribution under a Clayton [Clayton, 1978] copula model.
@@
Figure 1(a) shows how the independent censoring assumption fails to hold by omitting a confounding covariate – e.g.“tumor grade”. Since “tumor grade” relates to both $E$ (cancer relapse) and $C$ (cancer death), the variation in the covariate induces changes in both $E$ and $C$. That is, we want a model that predicts the time until relapse, but death is censoring. As “tumor grade” is predictive of tumor progression, an advanced (resp., early) stage is linked to shorter (longer) values for both cancer relapse $E$ and cancer death $C$, and consequently, $E$ and $C$ are not independent. This shows that the conditional independent censoring assumption might not hold if any confounding covariates are omitted or ignored in a regression model [Emura and Chen, 2018, Ch. 1]. This phenomenon is called residual dependencei.e.dependent censoring not adjusted by the covariates.
@@
Figure 1(b) demonstrates how dependent censoring biases the marginal survival estimates from the KM estimator [Kaplan and Meier, 1958]. Consider patient A, censored (due to cancer death) exactly one year after study entry, who likely would have experienced cancer relapse shortly thereafter. However, the KM estimator may overestimate this patient’s survival time by assuming that they are like those remaining in the study, who tend to live longer, because it does not consider that these censored patients tend to have shorter event times. Indeed, the estimator relies on a stepwise construction of the survival probability that depends on knowing how many individuals are at risk and the event probabilities at each time point [Bland and Altman, 1998].
To address such dependency, Figure 1(c) presents a copula-based modeling approach. A copula function $C_{\theta}(u,v)$ links random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1] and have recently gained interest in survival analysis [Foomani et al., 2023, Zhang et al., 2024, Liu et al., 2024]. Applied here, it captures the dependency between event and censoring distributions by sampling the marginals from a (survival) copula. In this example, we see lower tail dependence from the copula – describing the probability that small (early) values of one variable (e.g., event) are associated with small values of another variable (e.g., censoring).
In the absence of features, if any confounder related to the event censors the event time, we refer to this phenomenon as dependent censoring. However, it is typically impossible to verify this in practice because we only observe one outcome (either event or censoring) per instance, but not both [Tsiatis, 1975]. Although we can reduce dependent censoring simply by collecting more data (i.e., features) [Emura and Chen, 2018, Ch. 1], we would still have to evaluate our model under dependent censoring, independent of the learned model, preferably. Popular metrics for evaluating predicted survival probabilities (or risk scores) are Harrell’s concordance index (CI) [Harrell Jr et al., 1982], the Brier Score (BS) [Brier, 1950], and the mean absolute error (MAE) using one of hinge [Haider et al., 2020], margin [Haider et al., 2020] or pseudo times [Qi et al., 2023a]. Unfortunately, dependent censoring is both common [Emura and Chen, 2018, Ch. 1] and also introduces a bias that can affect the estimation of survival probabilities, and none of these metrics account for this. Alternative methods have been proposed to handle dependent censoring, such as Uno’s CI [Uno et al., 2011] using the inverse probability of censoring weighting (IPCW) [Robins and Finkelstein, 2000], but this approach still relies on the KM estimator. Motivated by these limitations, we introduce three new metrics for evaluating survival models under dependent censoring. To our knowledge, no prior work has proposed evaluation metrics that explicitly model the degree of dependency between $E$ and $C$. Our main contributions are:
  • We show that conventional survival metrics, such as Harrell’s CI, Uno’s CI, the IPCW-integrated Brier Score, and the marginal MAE, are biased under dependent censoring, meaning they can give inaccurate estimates of a model’s performance.
  • We propose three novel metrics – CI-Dep, IBS-Dep, and MAE-Dep – to evaluate survival models in the presence of dependent censoring. These metrics can evaluate the concordance (i.e., ranking), the accuracy of the survival curve, and the accuracy of the predicted time-to-event estimates, respectively, by considering the dependency between event and censoring.
  • We design a framework to create realistic, semi-synthetic datasets with various degrees of dependent censoring to determine if the proposed metrics can (1) report an unbiased error closest to the true error, and (2) rank the survival models by the true error.
  • We provide empirical evidence that our proposed metrics can provide more accurate error measures for survival models, mainly in terms of prediction accuracy.
Our code is publicly available at this GitHub repository.
@@ 2 Background and related work
2.1 Survival data and notation
Many survival analysis models estimate the probability that a specific event will occur for an individual at a time $T$ later than $t$, for all times $t>0$. Examples include predicting the time to death in patients with amyotrophic lateral sclerosis (ALS) [Kuan et al., 2023], predicting the time to a fall in older adults [Lillelund et al., 2024a], or evaluating a new experimental treatment for cancer [Zhang et al., 2011]. Most often, the observation for a given instance will consist of a random variable, $E$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $E$ can be characterized by the probability distribution function $F_{E}(t)=\text{Pr}(E\leq t)$ or, equivalently, by the survival function $S_{E}(t)=1-F_{E}(t)=\text{Pr}(E>t)$. Note that $S_{E}(t)$ (resp., $F_{E}(t)$) correspond to the probabilities of being event-free (resp., having experienced the event by) time $t$. We can estimate $S_{E}\left(t\right)$ nonparametrically, e.g., using the Kaplan-Meier [Kaplan and Meier, 1958] estimator (see Appendix H for a detailed derivation).
Definition 1
@@ 2.1 Survival data and notation
$\begin{split}\hat{S}(t)\ &=\prod_{t_{i}\leq t,d_{i}=1}\left(1-\frac{1}{n_{i}}\right),\quad\hbox{\rm for}\quad t\leq t_{\text{max}}\end{split}$ (2.1)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at-risk at time $t_{i}$; $\hat{S}(t)=1$ if no death occurs up to time $t$; $\hat{S}(t)$ is undefined for $t>t_{\text{max}}$ [Kaplan and Meier, 1958].
If features are entered into the model, popular methods include the CoxPH model [Cox, 1972, Katzman et al., 2018, Lillelund et al., 2024b], gradient-boosted regression [Ridgeway, 1999], Random Survival Forest [Ishwaran et al., 2008] and Multi-Task Logistic Regression (MTLR) [Yu et al., 2011]. In this context, we may define our survival dataset $\mathcal{D}$ as a set of observations, i.e., we have $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a feature (covariate) vector for instance $i$, $t_{i}\in\mathbb{R}^{\geq 0}$ is the time of either censoring or event, depending on which occurred first, and $\delta_{i}\in\{0,1\}$ is a event indicator where $\delta_{i}=0$ means the event has not happened by $t_{i}$ (right-censored) and $\delta_{i}=1$ means the event occurred at $t_{i}$ (uncensored). We consider each instance to have a potential event time $e_{i}$ and a right-censoring time $c_{i}$, where we observe only the earlier of the two. Therefore, we define the observed time and event indicator as:
          $t_{i}\ :=\ \min\{e_{i},c_{i}\},\quad\text{and}\quad\delta_{i}\ :=\ \mathbbm{1}[e_{i}\leq c_{i}],$
where $\mathbbm{1}[\cdot]$ represents the indicator function.
Given survival data $\mathcal{D}$, let $f_{E\mid\bm{X}}(\cdot)$ and $F_{E\mid\bm{X}}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon (i.e., from 00 to $t_{\text{max}}$). With these two functions, we then have the following useful definitions:
Definition 2
The survival function
$S_{E\mid\bm{X}}(t\mid\bm{X})=\text{Pr}(E>t\mid\bm{X})=1-F_{E\mid\bm{X}}(t\mid\bm{X}),$ (2.2)
represents the probability the event occurs at time $t$ [Gareth et al., 2021, Ch. 11].
Definition 3
The hazard function
$h_{E\mid\bm{X}}(t\mid\bm{X})$ =\ = $\lim_{\epsilon\to 0}\text{Pr}(E\in[t,t+\epsilon)\mid E\geq t,\bm{X})$ (2.3)
=\ = $\frac{f_{E\mid\bm{X}}(t\mid\bm{X})}{S_{E\mid\bm{X}}(t\mid\bm{X})}$
represents the instantaneous event rate at given time $t$, conditional on surviving $t$ [Gareth et al., 2021, Ch. 11].
@@ 2.2 Copula-Based Estimation
A copula is a function that links two or more random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. Among the various types of copulas, the Archimedean copula family is the most used, and enables modeling of dependencies in arbitrarily high dimensions using a single parameter, $\theta$. This family of copulas is defined by a generator function, $\varphi_{\theta}(t):[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3]. Here, we will focus on two common Archimedean copulas: Clayton [Clayton, 1978] and Frank [Frank, 1979].
Zheng and Klein [1995] generalized the KM estimator under independent censoring to the Copula-Graphic (CG) estimator under dependent censoring, which has become an important tool for analyzing survival data [Emura and Chen, 2016, Emura and Michimae, 2017, Hoora Moradian and Bellavance, 2019]. Rivest and Wells [2001] obtained the first explicit expression of the CG estimator under an assumed copula from the Archimedean family (e.g., Clayton or Frank). We can estimate the survival function using the CG estimator (see Appendix H for a detailed derivation):
Definition 4
@@
$\small\begin{split}\hat{S}_{E}(t)&=\varphi^{-1}\left(\sum_{{i:\,}t_{i}\leq t,\delta_{i}=1}\varphi\left(\frac{n_{i}-1}{n}\right)-\varphi\left(\frac{n_{i}}{n}\right)\right),\\ \hbox{for}&\quad t\ \leq\ t_{\text{max}}\end{split}$ (2.4)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at risk at time $t_{i}$; $\hat{S}_{E}(t)=1$ if no death occurs up to time $t$; $\hat{S}_{E}(t)$ is undefined for $t>t_{\text{max}}$ [Emura and Chen, 2018, Ch. 4].
Note that when $\varphi(\cdot)$ is uniform, this reduces to the standard KM estimator. See Appendix D for more details.

### v3

# Overcoming Dependent Censoring in the Evaluation of Survival Models

@@ Abstract
Conventional survival metrics, such as Harrell’s concordance index (CI) and the Brier Score, rely on the independent censoring assumption for valid inference with right-censored data. However, in the presence of so-called dependent censoring, where the probability of censoring is related to the event of interest, these metrics can give biased estimates of the underlying model error. In this paper, we introduce three new evaluation metrics for survival analysis based on Archimedean copulas that can account for dependent censoring. We also develop a framework to generate realistic, semi-synthetic datasets with dependent censoring to facilitate the evaluation of the metrics. Our experiments in synthetic and semi-synthetic data demonstrate that the proposed metrics can provide more accurate estimates of the model error than conventional metrics under dependent censoring.
@@ 1 Introduction
Learning survival prediction models resembles learning regression models, as both are essentially trying to learn, from a labeled dataset, a model that maps a description of an instance to a real value. Survival prediction examples include predicting the time to death in patients with amyotrophic lateral sclerosis (ALS) [1], predicting the time to a fall in older adults [2], or evaluating a new experimental treatment for cancer [3]. However, survival prediction differs from regression as some of the training instances are “censored” – that is, we only know a lower bound of the true time. This includes, for example, individuals who drop out or are still alive at the end of the study. The independent censoring assumption states that, within any subgroup of interest, censored patients at time $t$ are representative of all at-risk patients at that time with respect to their survival experience [4, Ch. 1]. Formally, we have $T\perp\mkern-9.5mu\perp C$, where $T$ is the event distribution and $C$ is the censoring distribution. This can also be stated explicitly in terms of the features $X$ as the conditional independent censoring assumption, i.e. $T\perp\mkern-9.5mu\perp C\mid X$ [5, Ch. 1] (see Appendix B for more details). Many standard survival tools – e.g., the Kaplan-Meier (KM) [6] and Nelson-Aalen [7, 8] estimators, and the Cox Proportional Hazards (CoxPH) model [9] – rely on the assumption of independent censoring [5, Ch. 1]. In the absence of features, if any confounder related to the event censors the event time, we refer to this phenomenon as dependent censoring.
@@
Figure 1: (a) Residual dependence between event time $T$ (cancer relapse) and censoring time $C$ (cancer death) denoted by the diamond “Dep”, due to an unobserved confounding covariate – here “Tumor grade”. (b) Evaluation under dependent censoring, where Patient A is censored immediately before the event for reasons related to the event. This leads to a biased estimation of the marginal survival distribution using the Kaplan-Meier (KM) [6] estimator. Note that the two curves coincide between the range $[0,c_{i}]$. (c) Scatter plot generated from a standard exponential distribution under an Archimedean Clayton [10] copula.
@@
Figure 1(a) shows an example of dependent censoring by omitting a confounding covariate – e.g.“tumor grade”. Since “tumor grade” relates to both $T$ (cancer relapse) and $C$ (cancer death), the variation in the covariate induces changes in both $T$ and $C$. That is, we want a model that predicts the time until relapse, but death is censoring. As “tumor grade” is predictive of tumor progression, an advanced (resp., early) stage is linked to shorter (longer) values for both cancer relapse $T$ and cancer death $C$, and consequently, $T$ and $C$ are not independent. This shows that the conditional independent censoring assumption might not hold if any confounding covariates are omitted or ignored in a regression model [4, Ch. 1].
@@
Figure 1(b) demonstrates how dependent censoring can bias the marginal survival estimates from the KM estimator [6]. Consider patient A, censored (due to cancer death) exactly one year after study entry, who likely would have experienced cancer relapse shortly thereafter. However, the KM estimator may overestimate this patient’s survival time by assuming that they are like those remaining in the study, who tend to live longer, because it does not consider that these censored patients tend to have shorter event times [11]. To overcome this limitation, we need a survival estimator that can account for dependent censoring when predicting the marginal distribution. To this end, Zheng and Klein [12] generalized the KM estimator to the Copula-Graphic (CG) estimator under an Archimedean copula, $C_{\theta}$, which leverages assumptions about the dependence structure between event and censoring. Figure 1(c) illustrates sampling from such a copula. In this example, we see lower tail dependence from the copula – describing the probability that small (early) values of one variable (e.g., event) are associated with small values of another variable (e.g., censoring). The CG estimator then uses this information when estimating the population-based survival probability.
To motivate our approach, it is important to distinguish model estimation and model evaluation. Under the (conditional) independence assumption, we can reduce dependent censoring at the estimation stage by simply collecting more data (i.e., features) and obtaining model parameters using e.g., standard CoxPH [9, 13, 14], Random Survival Forest (RSF) [15] or Multi-Task Logistic Regression (MTLR) [16]. Under the independence assumption, we can use recent copula-based techniques to jointly learn the parameters of the marginal distributions and the copula $C_{\theta}$ between them [17, 18, 19]. However, these approaches do not account for dependent censoring at the estimation stage, since standard survival metrics – such as Harrell’s concordance index (CI) [20], the Brier Score (BS) [21], and the mean absolute error (MAE) using one of hinge [22], margin [22] or pseudo times [23] – still assume that $T$ and $C$ are independent, regardless of the model parameters. If $T$ and $C$ are indeed dependent, assuming independence can lead to inaccurate performance estimates [4, Ch. 1].
Contribution.
To overcome these limitations, we introduce three new metrics – CI-Dep, IBS-Dep, and MAE-Dep – to evaluate survival models in the presence of dependent censoring. These new metrics can evaluate the concordance (i.e., ranking) of the predicted risk scores, the accuracy of the survival curve, and the accuracy of the time-to-event estimates, respectively. They achieve this by accounting for the dependency between event and censoring times using an Archimedean copula. We also design a framework to create realistic, semi-synthetic datasets, with dependent censoring to evaluate the existing and proposed metrics. We show empirically that our metrics can provide more accurate estimates of the model error than existing metrics, however, the performance ultimately depends on the dependence structure in the data and the censoring rate. To our knowledge, these are the first survival metrics that explicitly account for dependent censoring.
Our code is publicly available at: https://github.com/thecml/DependentEVAL/
@@ 2 Background and related work
2.1 Survival data
Many survival analysis models estimate the probability that a specific event will occur for an individual at a time $T$ later than $t$, for all times $t>0$. Most often, the observation for a given instance will consist of a random variable, $T$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $T$ can be characterized by the probability distribution function $F_{T}(t)=\text{Pr}(T\leq t)$ or, equivalently, by the survival function $S_{T}(t)=1-F_{T}(t)=\text{Pr}(T>t)$. Note that $S_{T}(t)$ (resp., $F_{T}(t)$) correspond to the probabilities of being event-free (resp., having experienced the event) by time $t$. We can estimate $S_{T}\left(t\right)$ nonparametrically, e.g., using the Kaplan-Meier [6] estimator (see Appendix G for a detailed derivation).
Definition 1.
@@ 2.1 Survival data
$\begin{split}\hat{S}_{T}(t)\ &=\prod_{t_{i}\leq t,d_{i}=1}\left(1-\frac{1}{n_{i}}\right),\quad\hbox{\rm for}\quad t\leq t_{\text{max}}\end{split}$ (2.1)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at-risk at time $t_{i}$; $\hat{S}(t)=1$ if no death occurs up to time $t$; $\hat{S}(t)$ is undefined for $t>t_{\text{max}}$ [6].
If features are entered into the model, we may define our survival dataset $\mathcal{D}$ as a set of observations, i.e., we have $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a feature (covariate) vector for instance $i$, $t_{i}\in\mathbb{R}^{\geq 0}$ is the time of either censoring or event, depending on which occurred first, and $\delta_{i}\in\{0,1\}$ is a event indicator where $\delta_{i}=0$ means the event has not happened by $t_{i}$ (right-censored) and $\delta_{i}=1$ means the event occurred at $t_{i}$ (uncensored). We consider each instance to have a potential event time $e_{i}$ and a right-censoring time $c_{i}$, and observe only the earlier of the two, so the observed time and event indicator are defined as $t_{i}:=\min\{e_{i},c_{i}\}$ and $\delta_{i}:=\mathbbm{1}[e_{i}\leq c_{i}]$, where $\mathbbm{1}[\cdot]$ denotes the indicator function. Given survival data $\mathcal{D}$, let $f_{T}=\text{Pr}(T=t)$ denote the density, and $S_{T}=\text{Pr}(T>t)$ the survival function, then we have the likelihood under (conditional) independent censoring:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[f_{T}(t_{i}\mid\bm{x}_{i})\,S_{C}(t_{i}\mid\bm{x}_{i})\right]^{\delta_{i}}\left[f_{C}(t_{i}\mid\bm{x}_{i})\,S_{T}(t_{i}\mid\bm{x}_{i})\right]^{1-\delta_{i}}.$ (2.2)
@@ 2.2 Copula-based estimation
A copula is a function that links two or more random variables by specifying their dependence structure [4, Ch. 1]. Among the various types of copulas, the Archimedean copula family is the most used, and specifies the dependence structure as a single parameter, $\theta$. This family of copulas is defined by a generator function, $\varphi_{\theta}(t):[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [4, Ch. 3]. Here, we will focus on two common Archimedean copulas: Clayton [10] and Frank [24]. Zheng and Klein [12] generalized the KM estimator (Equation 2.1) under independent censoring to the Copula-Graphic (CG) estimator under dependent censoring, which has become an important tool for analyzing survival data [25, 26, 27]. Rivest and Wells [28] obtained the first explicit expression of the CG estimator under an assumed copula from the Archimedean family (e.g., Clayton or Frank). We can estimate the survival function using the CG estimator (see Appendix G for a detailed derivation):
Definition 2.
@@
$\begin{split}\hat{S}_{T}(t)=\varphi^{-1}\left(\sum_{{i:\,}t_{i}\leq t,\delta_{i}=1}\varphi\left(\frac{n_{i}-1}{n}\right)-\varphi\left(\frac{n_{i}}{n}\right)\right),\quad\hbox{for}\quad t\ \leq\ t_{\text{max}}\end{split}$ (2.3)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at risk at time $t_{i}$; $\hat{S}_{T}(t)=1$ if no death occurs up to time $t$; $\hat{S}_{T}(t)$ is undefined for $t>t_{\text{max}}$ [4, Ch. 4]. Note that when $\varphi(\cdot)$ is uniform, this reduces to the standard KM estimator.
If features are entered into the model and given some Archimedean copula (e.g., Clayton or Frank), we can rewrite Equation 2.2 as the log-likelihood under dependent censoring (see Appendix E.3):
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta_{i}\;\text{log}\left[f_{T}(t_{i}|\bm{x}_{i})\right]\;+\;\delta_{i}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right]$
$+\;(1-\delta_{i})\;\text{log}\left[f_{C}(t_{i}|\bm{x}_{i})\right]+(1-\delta_{i})\;\text{log}\left[\frac{\partial}{\partial u_{2}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right]$ (2.4)
In this expression, the first term represents the log-likelihood of observing the event at time $t_{i}$. The second term captures the conditional probability of the censoring time occurring after the event time, given that the event occurred at $t_{i}$. By symmetry, the third and fourth terms mirror these quantities but with respect to the censoring time. Although Equation 2.2 may appear complex, the partial derivatives of any Archimedean copula have closed-form expressions, which allows us to compute the log-likelihood function using gradient descent [17].